Due to erratic sales of its sole product-a high-capacity battery for laptop computers-PEM, Inc., has been experiencing difficulty for some time. The companys contribution format income statement for the most recent month is given below: Sales (13,300 units x 40 per nit) $532,000 Variable expenses 319,200 Contribution margin Fixed expenses 212,800 236,800 Net operating loss S (24,000) Required: 1. Compute the companys CM ratio and its break-even point in both unit sales and dollar sales. CM ratio Break-even point in units Break-even point in dollars 2. The president believes that a $6,600 increase in the monthly advertising budget, combined with an intensified effort by the sales staff, will result in an $90,000 increase in monthly sales. If the president is right, what will be the effect on the companys monthly net operating income or loss? (Use the incremental approach in preparing your answer.) by
Answer #1

1. CM ratio = 212800*100/532000 = 40%

Break even dollar sales =Fixed cost/contribution margin ratio= 236800/.40 = 592000

Break even units sales = Break even dollar sales/sale price per unit = 592000/40 = 14800 units

c) Calculate Effect on net operating income

Sales (532000+90000) 622000
Variable cost (622000*60%) (373200)
Contribution margin 248800
FIxed cost (236800+6600) (243400)
Net income 5400

so net operating income will increase by (5400+24000) = 29400